Question 19 Consider two random variables X and Y with E(X)= 4, E(Y) = 2, E(XY) = 12, V(X) = 16 and V(Y) = 25, then the correlation coefficient between X and Y is: a. -0.2 b. -0.3 c. 0.2 d. 0.3 e. None of the above

need step by step distribution~
Answer #1

Correlation between x and y :

where

cov(x,y) = E(xy) - E(x) * E(y)

cov(x,y) = 12 - (4*2)

cov(x,y) = 12 - 8

cov(x,y) = 4

Correlation between x and y is 0.2